If $100 is invested at 4% compounded quarterly, what is the interest earned after:
3 years(interest earned)=
5 years(interest earned)=

Answer by Theo(13342)

f = p * (1+r)^n
f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

with quarterly compounding, the time period is 1 quarter.

.04 per year / 4 = .01 interest rate per quarter.
3 years * 4 = 12 quarters.
5 years * 4 = 20 quarters.

p = 100

f = p * (1+r)^n for p = 100 and r = .01 and n = 12 becomes:
f = 100 * (1.01)^12 which becomes:
f = 112.682503.

f = p * (1+r)^n for p = 100 and r = .01 and n = 20 becomes:
f = 100 * (1.01)^20 which becomes:
f = 122.019004